You work at a frozen yogurt shop during the summer. You need to order 5-ounce and 8-ounce cups. The storage room will only hold 10 more boxes. A box of 5-ounce cups costs $100 and a box of 8-ounce cups costs $150. A maximum of $1250 is budgeted for yogert cups. Write a system of linear inequalities that show the number of boxes of 5-ounce and 8-ounce cups that could be bought.
Answer by stanbon(75887)

You work at a frozen yogurt shop during the summer. You need to order 5-ounce and 8-ounce cups. The storage room will only hold 10 more boxes. A box of 5-ounce cups costs $100 and a box of 8-ounce cups costs $150. A maximum of $1250 is budgeted for yogert cups. Write a system of linear inequalities that show the number of boxes of 5-ounce and 8-ounce cups that could be bought.
------------
Quantity Equation: f + e = 10
Value Equation...:100f + 150e = 1250
----------------------------------------
Modify:
100f + 100e = 1000
100f + 150e = 1250
----------------------------
Subtract the 1st from the 2nd to solve for e:
50e = 250
e = 5 boxes (# of boxes of eight-once cups)
f + e = 10, so f = 5 boxes: (# of boxes of five-once cups)
-----------------------------------------------------------------
Cheers,
Stan H.